# Tianhao Wei∗,1, Liqian Ma∗,1,2, Rui Chen1, Weiye Zhao1, Changliu Liu1

### v1

@@ 4 Experiment
Figure 4: Three manipulation tasks that have inherently different challenges and requirements. The balance task requires an accurate, high-frequency, feedback controller. The open door task requires properly handling articulated objects, the executed trajectory has to perfectly match the swing path. The safe pick and place task requires guaranteeing collision avoidance for the whole robot arm.
@@
The experiment is designed to manifest the following features: 1) Meta-Control enables synthesizing skills for heterogeneous requirements; 2) Meta-Control exploits control knowledge from LLM; 3) The Meta-Control pipeline is critical to the success; 4) Meta-Control works on a wide range of open world tasks; 5) Meta-Control enables model-based analysis, providing formal guarantees on various properties. 6) Meta-Control transfers to real robots and different embodiments easily.
We implement our pipeline with Drake[76], a framework designed for model-based control. For the hardware experiment, we used a Kinova Gen3 robot arm. The language model is GPT 4.0.
@@ 4.1 Meta control enables synthesizing challenging skills
As shown in fig. 3 and fig. 4, Meta-Control successfully synthesizes controllers for various challenging tasks with inherently different requirements, both in simulation and the real world. The task challenges and synthesized control systems are summarized in table 1. A more detailed description can be found in appendix B. fig. 5 shows that Meta-Control can identify core metrics and critical parameters that affect performance, then efficiently and effectively tune the parameters based on the metric through trial and error. The complete process of skill synthesis can be found at appendix C.
Task Challenge Meta-Control designed system
Figure 6: Meta-Control generalizes to different embodiments because the synthesized controller is fully model-based. A controller synthesized on Kinova can transfer to Franka Pranda simply by replacing the robot dynamic model.
Method API API + Formulation API + Template Meta-Control
@@
Table 2: The ablation study shows that hierarchical formulation (Formulation) and simplification by dynamic model and controller templates (Templates) both improves the success rate, compared to directly synthesis with only the low-level API.
4.2 Meta control exploits dynamics priors
In contrast to previous work that focuses more on spatial relationship priors [1, 4, 5], Meta-Control exploits the LLM’s internalized knowledge of dynamics, enabling to synthesize high-performance controllers rigorously. In the balance cart pole task, LLM designs the task space dynamics $h(z,v)$ with a linear approximation in the form of $\dot{z}=Az+Bv$ around the upright position of the pole, where $A$ and $B$ are directly given by the LLM: $A=[[0,1,0,0];[0,0,\frac{m_{\text{pole}}g}{m_{\text{cart}}},0];[0,0,0,1];[0,0,\frac{g(m_{\text{cart}}+m_{\text{pole}})}{l_{\text{pole}}m_{\text{cart}}},0]]$, $B=[0;\frac{1}{m_{\text{cart}}};0;-\frac{1}{l_{\text{pole}}m_{\text{cart}}}]$, where $m$ denotes mass, $l$ denotes length, $g$ is the acceleration of gravity. Detailed description of the synthesized control system can be found at appendix B.
@@ 4.3 Ablation study of the Meta Control pipeline
To demonstrate the necessity of the hierarchy formulation and simplification using templates, we show the success rate of control system synthesis under different conditions. We test the success rate in 3 stages: design (via templates), implementation (interface alignment), and execution. We say that a design is successful if the LLM-designed control system has the potential to perform the skill judged by an expert. We say that the implementation is successful if the system can run without errors. We say that the execution is successful if the control system finishes the skill as desired. We repeat each task $10$ times to compute the success rate. Randomness is caused by the LLM with a default temperature $1.0$. As shown in table 2, we can see that the success rates of all steps in the baseline are lower than Meta-Control. Although LLM gives reasonable architectures to finish the task in the Balance and Open Door task, the LLM fails to provide correct code and parameters to realize the control system due to the complexity of the system and the huge design space. With all the modules, we achieve the highest success rate for all tasks. The Safe Pick&Place task is especially difficult for the baseline because the baseline methods, even though prompted to avoid collision, were unable to successfully design a controller that can avoid collision continuously.
4.4 Generalization to different attributes/states
Meta-Control synthesized control systems can easily generalize to scenarios of different attributes/states due to the model-based nature. Given a successfully synthesized control system, we test different attributes/states and calculate the success rate. The range of change and the results are shown in table 3. Examples are shown in fig. 7. Meta-Control achieved a 100% success rate for all scenarios.
4.5 Meta Control enables Model-based Analysis
One merit of our method is that the synthesized controller is fully model-based, allowing a rigorous formal analysis for a variety of properties:
Convergence and Stability: For the balance cart-pole task, we can provide a guarantee of convergence by solving the Riccati equation for the LQR controller. The closed-loop system matrix $A-BK$ has the following four eigenvalues: $-412.29,-9.925,-1.502+1.175j,-1.502-1.175j.$ All of them have negative real parts, which means that the system is guaranteed to converge. More rigorous analysis can be conducted by taking the linearization error into consideration.
4.6 Transfer to real robot and different embodiments
The control system for opening doors is synthesized in simulation and is executed in both simulation and the real world. As shown in fig. 3 and fig. 4, the behavior is consistent and no sim-to-real gap is observed because the synthesized controller is model-based, closed-loop, highly explainable, and math-guaranteed. The controller can also generalize to different embodiments with the same low-level API (e.g. 7 DoF joint torque) given the model of the new embodiments, shown in fig. 7.
@@ 4.6 Transfer to real robot and different embodiments
Table 3: The synthesized controllers easily generalize to scenarios with different object states/attributes. The left column lists the range of the parameters for each scenario. The right column indicates the success rate (out of 10 trials) for each set of parameters.
@@ 5 Limitation and Discussion
In this work, we propose Meta-Control, a novel framework for zero-shot model-based control system synthesis using LLM, tailored for heterogeneous robotic tasks. Through both simulations and real-world tests, we demonstrate Meta-Control’s potential to expand the autonomous capabilities of robots. Despite compelling results, Meta-Control has several limitations. Such as dependency on accurate system state estimation, which may not be available in some open-world tasks, and dependency on predefined dynamic models and control templates, which may restrict adaptability to completely new heterogeneous tasks. The synthesis process also demands substantial computational resources, which could hinder real-time synthesis. Future directions include enabling automatic perception selection, incorporating learning-based templates, and accelerating synthesis speed.
@@ 6 Acknowledgement
Appendix A Joint torques during openning a door
Appendix B Task descriptions
@@ Appendix C Full conversation
The full conversation on skill synthesis is attached below. Some long numerical arrays are omitted for the sake of clarity. Controller templates, dynamical model templates, and input port samples are attached after the conversation.
C.1 Design via Templates

### v3

@@ 4 Experiment
Figure 4: Five manipulation tasks that have inherently different challenges and requirements. For example, the balance task requires an accurate and high-frequency feedback controller. The safe pick and place task requires guaranteeing collision avoidance for the whole robot arm. The open door task requires properly handling articulated objects; and the executed trajectory has to perfectly match the swing path.
@@
The experiment is designed to manifest the following features. We first show that Meta-Control enables synthesizing skills for heterogeneous requirements by exploiting the control knowledge of LLM. Then we did an ablation study to show that all three levels of the Meta-Control pipeline improve the success rate; Finally, we reveal the benefit of the model-based design, including robustness, generalizability, and formal guarantees.
Most of the experiments are conducted on four representative tasks. These tasks cover constraints that appear frequently in the real world, such as constraints on position, velocity, force, compliance, convergence, and safety. We implement our pipeline with Drake[74], a framework designed for model-based control. For the hardware experiment, we used a Kinova Gen3 robot arm. The language model is GPT 4.0 with a default temperature $1.0$.
@@ 4.1 Meta control enables synthesizing challenging skills
As shown in fig. 2 and fig. 4, Meta-Control successfully synthesizes controllers for various challenging tasks with inherently different requirements, both in simulation and in the real world. Figure 5 shows that Meta-Control can identify core metrics and critical parameters that affect performance, then efficiently and effectively tune the parameters based on the metric through trial and error. A detailed description of task challenges and synthesized control systems can be found in section A.1. The complete skill synthesis process can be found in appendix C.
Meta-Control exploits dynamics priors In contrast to previous work that focuses more on spatial relationship priors [1, 4, 5], Meta-Control exploits the LLM’s internalized knowledge of dynamics, enabling rigorous synthesis of high-performance controllers. In the balance cart pole task, LLM designs the task space dynamics $h(z,v)$ with a linear approximation in the form of $\dot{z}=Az+Bv$ around the upright position of the pole, where $A$ and $B$ are correctly given by the LLM: $A=[[0,1,0,0];[0,0,\frac{m_{\text{pole}}g}{m_{\text{cart}}},0];[0,0,0,1];[0,0,\frac{g(m_{\text{cart}}+m_{\text{pole}})}{l_{\text{pole}}m_{\text{cart}}},0]]$, $B=[0;\frac{1}{m_{\text{cart}}};0;-\frac{1}{l_{\text{pole}}m_{\text{cart}}}]$, where $m$ denotes mass, $l$ denotes length, $g$ is the acceleration of gravity. Detailed description of the synthesized control system can be found at section A.1.
Method API API + Hierarchy API + Template Meta-Control
@@
Table 1: The ablation study shows that hierarchical formulation (Hierarchy) and templates-based synthesis (Templates) both improves the success rate, compared to directly synthesis with only the low-level API. API + Template can be viewed as a generalized motion-primitive method.
@@ 4.2 Ablation study of the Meta Control pipeline
To demonstrate the necessity of the hierarchy formulation and simplification using templates, we show the success rate of control system synthesis with different ways to query the LLM. Meta-Control use the query: “output is 7-DoF torque, input is …” (API) + “consider a task space and a tracking space…” (Hierarchy) + “use available templates” (Template). We test the success rate in 3 stages: design (via templates), composition (interface alignment), and execution (after few-shot optimization). We say that a design is successful if the LLM-designed control system has the potential to perform the skill judged by an expert. We say that the implementation is successful if the system can run without errors. We say that the execution is successful if the control system finishes the skill as desired. We repeat each task $10$ times to compute the success rate. Randomness is caused by the LLM. As shown in table 1, we can see that the success rates of all steps in the baseline are lower than Meta-Control. Although LLM gives reasonable architectures to finish the task in the Balance and Open Door task, the LLM fails to provide correct code and parameters to realize the control system due to the complexity of the system and the huge design space. With all the modules, we achieve the highest success rate for all tasks. The Safe Pick&Place task is especially difficult for the baseline because the baseline methods, even though prompted to avoid collision, were unable to successfully design a controller that can avoid collision continuously.
4.3 Meta-Control brings benefits of model-based design
The model-based nature of the synthesized controller brings a variety of benefits, such as robustness, explainability, generalizability, and rigorous analysis.
Generalization to different attributes/states Meta-Control synthesized control systems can easily generalize to scenarios of different attributes/states due to the model-based nature. Given a successfully synthesized control system, we test different attributes/states and calculate the success rate. The range of change and the results are shown in table 2. Examples are shown in fig. 7. Meta-Control achieved a 100% success rate for all scenarios.
@@ 4.3 Meta-Control brings benefits of model-based design
Table 2: The synthesized controllers easily generalize to scenarios with different object states/attributes. The left column lists the range of the parameters for each scenario. The right column indicates the success count of 10 trials for each set of parameters.
Rigorous Formal Analysis The model-based design allows rigorous formal analysis for a variety of properties, making autonomous certification by LLM possible. For example, we can guarantee the convergence for the cart-pole stask by solving the Riccati equation, and ensure collision avoidance for the pick-and-place task by certifying a safety index. Details can be found in section A.3
Transfer to real robot and different embodiments The control system for opening door is synthesized in simulation and is executed both in simulation and in the real world. As shown in fig. 2 and fig. 4, the behavior is consistent and no sim-to-real gap is observed because the synthesized controller is model-based, closed-loop, highly explainable, and math-certified. The controller can also generalize to different embodiments with the same low-level API (e.g. 7 DoF joint torque) given the model of the new embodiments as shown in fig. 7.
Figure 7: Meta-Control generalizes to different embodiments because the synthesis is fully model-based. Furthermore, a controller synthesized on Kinova can transfer to Franka Pranda simply by replacing the robot dynamic model thanks to the model-based nature.
@@ 5 Limitation and Discussion
In this work, we present Meta-Control, a novel framework for automatic model-based control system synthesis using LLMs, designed for heterogeneous robotic tasks. Simulations and real-world tests demonstrate its potential to enhance robotic autonomy. However, Meta-Control has limitations, such as reliance on accurate system state estimation and predefined models, which limit adaptability to new tasks. The LLM can also make suboptimal decisions, like selecting PID over LQR, complicating parameter tuning. Future improvements include better reflection across the pipeline, stronger reasoning capabilities, faster synthesis, and the integration of learning-based templates and automatic perception selection.
@@ 6 Acknowledgement
Appendix A Experiment details
A.1 Task descriptions
Task Challenge Meta-Control designed system
@@ A.1 Task descriptions
A.2 Joint torques during openning a door
A.3 Formal Analysis of Properties
Convergence and Stability: For the balance cart-pole task, the convergence can be guaranteed by solving the Riccati equation for the LQR controller. The closed-loop system matrix $A-BK$ has the following four eigenvalues: $-412.29,-9.925,-1.502+1.175j,-1.502-1.175j.$ All of them have negative real parts, which means that the system is guaranteed to converge. More rigorous analysis can be conducted by taking the linearization error into account.
A.4 Runtime and parameter optimization rounds
In our experiments, it usually takes GPT-4o 1-3 minutes to synthesize a skill, depending on the capacity of the servery.
Appendix B Prompt Design
We use a variety of techniques to improve the quality of the full dialogue. First, we introduce some techniques in the overall design, followed by the techniques used in each step of the three-level pipeline.
B.1 Overall Design Techniques
  1. 1.
    Divide skill synthesis into three dialogues, corresponding to the three-level pipeline. This approach avoids overly long contexts, enables step-by-step reasoning, and allows reflections to focus on the current step.
  2. 2.
    At the end of each dialogue, Meta-Control asks the LLM to generate a summary to warm-start the next dialogue, including the decisions made in the current step and all necessary information for the next step.
  3. 3.
    Use a generic prompt with placeholder design to make the prompt zero-shot generalizable while remaining task-specific. Each placeholder is uniquely identified by a <step_name>. The placeholders are dynamically replaced by the LLM generated code.
  4. 4.
    Provide guidance and examples to help the LLM generate code following a specific format. The code should start with a <step_name>, e.g., <task_model>, followed by a Python code block encompassed by triple backticks (‘‘‘).
  5. 5.
    Use <step_name> to insert LLM-generated code into the control system. Our code extractor uses regular expressions to search for <step_name>, extract code from the LLM output, and then insert the code into the control system based on <step_name>.
  6. 6.
    Use a closed-loop design procedure. The LLM-generated code is executed by our program, which receives feedback, such as error messages or measurements during execution. Meta-Control then asks the LLM to improve the code based on the feedback.
  7. 7.
    Use code replacement instead of code completion to progressively construct the control system. Debugging and improvement are difficult with an incomplete control system code that is only executable after completing it with LLM-generated code, as feedback for intermediate steps would be unavailable. Therefore, we initially provide a blank but executable control system code, which outputs all zero joint torques. At each step of the synthesis, part of the code is replaced based on the identifier. We designed the control system to be always executable during the design process as long as the generated code is correct.
B.1.1 Dialogue 1: Design via Template Prompt
  1. 1.
    Use a role-play prompt (e.g., a professor in control and a proficient programmer) to invoke control knowledge and coding skills.
  2. 2.
    Provide some optional principles to assist the design, such as “design the task controller on simplified dynamics” and “use the task controller to accomplish the primary objective while using the tracking controller to satisfy constraints or secondary objectives.”
  3. 3.
    Use model templates and controller templates to ground the design space. The templates greatly reduce the design space and anchor the behavior of the synthesized control system, making it explainable and provable.
  4. 4.
    Provide all available measurements and their samples. The sample measurements give the LLM a precise idea of what inputs are available and their format, helping it to decide on the appropriate task controller and tracking controller.
  5. 5.
    Provide an example of the code that the LLM should return, including the <step_name> and how to initialize the chosen models and controllers.
  6. 6.
    Provide a checklist of requirements that the LLM should satisfy to reduce errors, such as “choices have been made for all steps” and “<step_name> must be included before each code block.”
  7. 7.
    Use reflections. Meta-Control extracts and evaluates the code from the LLM response, then informs the LLM if the initialization of the chosen models and controllers causes any errors, allowing the LLM to decide if any fixes or improvements are needed. This process is repeated until there are no errors or the maximum iteration is reached.
  8. 8.
    Ask the LLM to generate a summary of the design to warm-start the composition dialogue. The summary includes the chosen models and controllers, a detailed description of the inputs for the controllers, the meaning of each dimension, and any other necessary information.
B.1.2 Dialogue 2: Composition via Semantic Understanding
  1. 1.
    Use a role-play prompt as in the previous dialogue.
  2. 2.
    Use the design summary to warm-start the dialogue, providing the task, the chosen controllers and models, their input and output ports, and how they should be combined into a control system to complete the task.
  3. 3.
    Provide examples of how to access the values of the input and output ports of a model or controller.
  4. 4.
    Provide an example of the input converter, demonstrating how to connect the input and pad zeros when necessary. The input converter is a set of callback functions that assign values to the input based on the output of other modules.
  5. 5.
    Provide tips and potential errors to avoid when implementing the input converter. For example, “Ensure the dimensions of the output match the port. The output value may only be partially available from the input port,” and “For pose, velocity, and force vectors, rotation always comes first.”
  6. 6.
    Use partial updates. Because input converters involve many functions, if an error is detected during reflection, Meta-Control only updates the function that the LLM identifies as causing the error, instead of updating all functions, which may cause more errors if the code is not exactly the same.
  7. 7.
    Ask the LLM to generate a summary of the data flow to warm-start the parameter tuning, including parameters of the controller, parameters of the model, input converter structures, and potential performance metrics.
B.1.3 Dialogue 3: Parameter Optimization
  1. 1.
    Use a role-play prompt as in the previous dialogues.
  2. 2.
    Provide a list of code blocks that the LLM can update to optimize the parameters, covering models, controllers, and input converters.
  3. 3.
    Warm-start the dialogue with the data flow summary, providing a detailed implementation of the control system, a description, and samples of the module inputs.
  4. 4.
    Provide a guide for parameter optimization: identify key parameters, identify performance metrics from task state and task control, and then determine if the time series of the metrics is desired.
  5. 5.
    Provide a checklist of errors to identify the cause of degraded performance, including definition, initialization, implementation, and suboptimal parameters.
  6. 6.
    Iteratively optimize the parameters with reflections. Every time the LLM updates the code or parameters, Meta-Control evaluates the new system, provides a time series of the chosen metrics, and allows the LLM to decide if the performance is desired. This process is repeated until the performance is satisfactory.
@@ Appendix C Full conversation
The full conversation on skill synthesis is attached below. Some long numerical arrays are omitted for the sake of clarity. The controller templates, the dynamical model templates, and the input port samples are attached after the conversation.
C.1 Balance
C.1.1 Design via Templates
@@ C.1.3 Parameter Optimization
C.2 Pick and Place
C.2.1 Design via Templates
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]markdownprompt/conversation/pick_and_place/pick_and_place_design.md
C.2.2 Composition via Semantic Understanding
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]markdownprompt/conversation/pick_and_place/pick_and_place_implement.md
C.3 Open Door
C.3.1 Design via Templates
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]markdownprompt/conversation/open_door/open_door_design.md
C.3.2 Composition via Semantic Understanding
[…]
[fontsize=, breaklines, frame=lines, framesep=2mm]markdownprompt/conversation/open_door/open_door_implement.md